# Branches, Assemble!  Multi-Branch Cooperation Network for Large-Scale Click-Through Rate Prediction at Taobao

### v1

@@ 3.1. Overview
In CTR prediction, given the raw input feature $\boldsymbol{x}=\{\boldsymbol{x_{1}},\boldsymbol{x_{2}},...,\boldsymbol{x_{M}}\}$ with $M$ feature fields and user feedback label $\boldsymbol{y}\in\{0,1\}$, we aim to learn a mapping function $\mathcal{F}:\boldsymbol{x}\rightarrow\boldsymbol{y}$ to predict user behaviors for online services. Usually, $\mathcal{F}$ denotes different deep neural networks (e.g., DCNv2 (Wang et al., 2021)) and $\boldsymbol{x_{i}}$ denotes the $i$-th feature field that can be categorical, multi-valued, or numerical feature of users or items. Especially in industrial scenarios, the number of feature fields $M$ can be in hundreds or even thousands.
@@
3.2. Expert-based Feature Grouping and Crossing (EFGC)
The EFGC branch is designed to improve the model’s memorization capability for specific feature pairs, guided by domain-expert knowledge. It consists of two main components: the embedding layer, and the feature grouping and crossing module.
@@ 3.2. Expert-based Feature Grouping and Crossing (EFGC)
Feature Grouping: In CTR prediction, different feature fields have different physical meanings, and their combinations are crucial for personalized ranking. For instance, in e-commence search, combining user profile feature fields (e.g., userID, age) and item profile feature fields (e.g., itemID, price, sales volume) can reveal a user’s general preferences regarding item price and quality. Conversely, combinations like item title and item category could show less contributions to personalized ranking. In essence, different feature field combinations have varying degrees of influence on CTR prediction. The common practice of concatenating $\boldsymbol{e}_{i}$ in Eq. 1 may fail to highlight certain specific combinations and might introduce irrelevant feature patterns into subsequent deep modules.
Table 1. The designed feature groups in our image2product search.
group feature fields intention
@@
It is crucial to organize feature fields into distinct groups and perform feature crossing within these specific groups, driven by domain-expert knowledge. To achieve this, we can follow three main steps: 1) Identify Desired Matching Patterns: Determine the types of matching patterns that need to be emphasized based on the problem domain. 2) Group Feature Fields Accordingly: Select and group the feature fields in a way that aligns with the identified matching patterns. 3) Integrate Group Embeddings into Subsequent Layers: Input the embeddings from each group into the following layers to facilitate the learning of feature interactions. For instance, in step 1 of our image2product search, we aim to emphasize the visual relevance between query image and item image, while also personalizing the results based on each user’s preferences for specific item attributes. Step 2 is formulated as:
@@
Remark: Through splitting feature fields into groups and performing feature crossing within these groups, EFGC branch enhances the model’s ability to memorize specific patterns and avoid undesired feature crossing. Additionally, it is flexible and can be extended to include more types of feature groups. We note feature engineering techniques also use specific feature fields to create new features, such as the count of clicks between userID and brand. Both feature engineering and EFGC employ domain-expert knowledge. However, the difference is feature engineering creates raw input features, whereas EFGC works in a learning-based manner, enabling the automatic discovery of patterns among feature fields within groups. They are not mutually exclusive and can be used together in practice.
@@ 3.3. Deep Net and Low Rank Cross Net
where $\mathbb{FC}$ denotes one MLP layer for dimension reduction. $f^{\text{Cross}}$ denotes a mixture of low-rank cross network blocks. Due to the page limit, we do not elaborate it in detail here. For a detailed formulation, please refer to (Wang et al., 2021). After the feature interaction learning of each branch, we further include a shared top layer $f_{\theta}(\cdot)$ to reduce parameters and regularize training. The output latent features of different branches are formulated as:
@@
3.4. Multi-branch Cooperation Scheme
@@ 3.4.1. Branch Co-teaching
Different branches of MBCnet possess distinct advantages and inherently focus on modeling various patterns. These patterns, in turn, are represented by the training samples. As a result, the learning capabilities of different branches can vary significantly across different samples. On particular samples, if a robust learning branch has been effectively trained, it can assist a comparatively weak learning branch in enhancing its performance. We formulate this collaboration idea as ”branch co-teaching” objective function, whereby given specific samples, the strong branch provides guidance to the weak branch in training.
@@
where $BCE$ means binary cross entropy loss. Since CTR prediction is usually regarded as a binary classification task, we employ $-\log(0.5)$ as a threshold to assess the learning progress of each branch [^3] The prediction probabilities may be affected by the data distribution, we experimented with various thresholds and found $-\log(0.5)$ generally did the best.. When the loss value is less than $-\log(0.5)$, we consider the branch on these samples is learned effectively. Otherwise, we consider the branch is not learned well.
Branch Co-teaching Objective Formulation: Given the above defined strong and weak branches, we can write the co-teaching loss as follows:
@@
where $i,j\in\{\text{EFGC},\text{Deep},\text{Cross}\}$. SG means stop gradient. $I^{ij}\in\{0,1\}$ is the indicator that only equals 1 when $\boldsymbol{p^{i}}$ is strong and $\boldsymbol{p^{j}}$ is weak shown in Eq. 11. Similarly, $I^{ji}\in\{0,1\}$ is the indicator that equals 1 only when $\boldsymbol{p^{j}}$ is strong and $\boldsymbol{p^{i}}$ is weak shown in Eq. 10. $C$ is the number of non-zero values calculated from $I^{ij}$ and $I^{ji}$, indicating how many disagreement pairs are used in $\mathcal{L}_{BCT}$. In this equation, we take the prediction of strong branches as soft label to guide the learning of weak branches on selected samples.
Remark: The main idea of this branch co-teaching objective is that the teaching guidance is only conducted on selected samples where two branches show disagreed predictions. It is one important difference between this approach and the knowledge distillation in (Zhu et al., 2020) which performs full sample supervision between teachers and students. Given specific samples, learning abilities of different branches vary and are gradually improved, which suggests that the branches cannot offer appropriate supervision for all samples, especially in the early training stage. Employing supervision on all samples could easily introduce noisy supervision, and this is why EnKD (Zhu et al., 2020) uses pre-trained teachers and temperature scaling tricks to mitigate the effects of such noise. In contrast, our branch co-teaching operates in an end-to-end manner, focusing solely on samples with disagreeing predictions to minimize supervision noise.
@@ 3.4.2. Moderate Differentiation
In MBCnet, multiple feature interaction branches make predictions for the same input sample. When these branches are supervised on the same samples, they have risk to learn identical feature interaction patterns. This redundancy can significantly hinder the model’s capacity to explore a variety of feature interactions. To enhance the model’s robustness, it is better to enable distinct branches to learn differentiated latent features, encouraging them to capture diverse interaction patterns even on the same sample. However, excessive differentiation between these latent features may disrupt the consistency of shared patterns across branches, potentially harming model performance. In this context, we propose ”Moderate Differentiation” for the latent features $\boldsymbol{z^{i}}$ across branches. This approach promotes the exploration of diverse interaction patterns while ensuring a balance between differentiation and consistency.
@@
where $W^{ji},W^{ij}\in\mathbb{R}^{d\times d}$ are two learnable matrices and they are inverses of each other. To avoid feature collapse where multiple $\boldsymbol{z^{i}}$ collapse to one point after transformation, we employ orthogonal transformation, a specification of equivalent transformation. This means $W^{ji}$ and $W^{ij}$ are the transpose to each other, i.e., $W^{ij}=(W^{ji})^{T}$. The orthogonal transformation has an advantage of preserving inner product of vectors, namely it can keep similarities of vectors after transformation.
Moderate Differentiation Regularization: Taking the above transformation constraint analysis into consideration, we can write the regularization loss for moderate differentiation of branches as:
@@ 3.4.3. Branch Fusion and Training Objective Function
where $\alpha$ and $\beta$ are two hyper-parameters to weight the importance of loss terms. Notice that we use an arbitrary sample to succinctly demonstrate above objective function with minimal symbols. When using the objective function in practice, we compute the expectation of loss values of mini-batch samples during training.
Discussion: Popular CTR works (Cheng et al., 2016; Wang et al., 2021; Guo et al., 2017; Mao et al., 2023) investigate various feature interaction techniques for improved performance. However, each technique offers its own advantages, and relying on just one may not adequately capture the patterns in complex datasets, especially in industrial scenarios with enormous users and items. Our MBCnet incorporates a novel cooperation mechanism for different feature interaction approaches to better exert their learning power. Compared to recent research (Zhu et al., 2020) that also leverages supervision from other models, MBCnet implements a distinct and more advanced learning framework. The comparative results presented in the following sections will highlight this difference.
@@ 4.1.1. Datasets
We conduct experiments on two industrial datasets: Pailitao-12month and Pailitao-24month. Both are collected from ourimage2product retrieval, also known as “Pailitao” at Taobao. Pailitao-12month and Pailitao-24month contain user search behaviors ranging from 2023-06-01 to 2024-05-31 and 2022-06-01 to 2024-05-31, respectively. For both datasets, there are various feature fields, including query image feature, user profiles, user statistical features, item image feature, item attributes, item statistical features and some other designed clustering and crossing features with expert knowledge. All experimental models utilize the same input features to ensure a fair comparison of their modeling differences. The statistics of datasets are provided in Table 2.
@@ 4.1.2. Baselines
We leverage different strong algorithms for comparison including single branch methods and branch ensemble methods. The single branch methods are as follows. 1) DNN is a deep multi-layer perception model, 2) CrossNetV2 (Wang et al., 2021) is a low-rank version of cross net (Wang et al., 2017a) to explicitly learn high-order feature interactions. The branch ensemble methods are listed as follows. 3) Wide&Deep (Cheng et al., 2016) combines a wide set of cross-product feature transformations and deep neural networks. 4) DCNv2 (Wang et al., 2021) contains mixture-of-expert based low-rank feature crossing and DNN to learn bounded-degree feature interactions. 5) MMOE+ (Ma et al., 2018) originally is a multi-task learning model. We employ its mixture-of-expert encoder here to work as a baseline with multiple-expert ensemble. 6) EnKD (Zhu et al., 2020) studies knowledge distillation technique for model ensemble. 7) FinalMLP (Mao et al., 2023) designs feature gating and interaction aggregation layers to make an enhanced two-stream MLP model.
Table 3. The AUC and LogLoss comparison results of different models on our two datasets. Best results are in bold and the most competitive baseline results are underlined.
@@ 4.1.3. Parameter Settings
In our experiments, we split the datasets by chronological order, with the last seven-day data as validation and test set, and the rest is taken as the train set. For all methods, we set the latent dimension of last feature learning layer as 128. The validation performance is set as early-stop condition in training. In MBCnet, $\alpha$ and $\beta$ both equal 0.1 by hyper-parameter searching. In EFGC branch, hidden units of each group crossing are [1024,128]. In CrossNet branch, the number of experts is 2 and each expert has 2 layers with low rank dimension as 16. The hidden units of dimension reduction layer in EFGC and CrossNet are both 512. In Deep branch, the hidden units are [2048,1024,512,512,512]. Hidden units of the shared top layer are [512,256,128]. Hyper-parameters of baselines are searched to ensure each achieves its best performance. Loss codes will be made public later.
@@ 4.2.1. Offline Comparison
We make model comparisons with recent competitive baselines on our two large-scale industrial datasets. The results are given in Table 3. From this table, we can observe that: 1) our MBCnet achieves obvious improvement over recent competitive baselines. It has a 0.61% and 0.72% AUC increase over the most competitive model on two datasets, which are impressive improvements in large-scale industrial data scenarios. 2) On Pailitao-24month dataset, single DNN outperforms other baselines. This trend is commonly observed in large-scale industrial settings, where the enormous data can overwhelm many modeling techniques. Even in this context, MBCnet still achieves superior improvements, demonstrating the effectiveness of our approach that leverages cooperative multi-branches to predict user behaviors.

### v2

@@ 3.1. Overview
In CTR prediction, given the raw input feature $\boldsymbol{x}=\{\boldsymbol{x_{1}},\boldsymbol{x_{2}},...,\boldsymbol{x_{M}}\}$ with $M$ feature fields and user feedback label $\boldsymbol{y}\in\{0,1\}$, we aim to learn a mapping function $\mathcal{F}:\boldsymbol{x}\rightarrow\boldsymbol{y}$ to predict user behaviors for online services. Usually, $\mathcal{F}$ denotes different deep neural networks (e.g., DCNv2 (Wang et al., 2021)) and $\boldsymbol{x_{i}}$ denotes the $i$-th feature field that can be categorical, multi-valued, or numerical feature of users or items. Especially in industrial scenarios, the number of feature fields $M$ can be in hundreds or even thousands, bringing great challenges of modeling feature interactions.
@@
3.2. Extensible Feature Grouping and Crossing
The EFGC branch is designed to improve the model’s memorization capability for intended interactions of specific feature fields, guided by domain-driven knowledge. It consists of two main components: the embedding layer, and the feature grouping and crossing module.
@@ 3.2. Extensible Feature Grouping and Crossing
Feature Grouping: In CTR prediction, different feature fields have different physical meanings, and their interactions are crucial for personalized ranking. For instance, in e-commence search, combining user profile feature fields (e.g., userID, age) and item profile feature fields (e.g., itemID, price, sales volume) can reveal a user’s general preferences regarding item price and quality. Conversely, interactions like item title and item category could show less contributions to personalized ranking. In essence, different feature field interactions have varying degrees of influence on CTR prediction. The common practice of concatenating $\boldsymbol{e}_{i}$ in Eq. 1 may fail to highlight certain specific interactions and introduce irrelevant feature patterns into subsequent deep modules.
Table 1. The designed feature groups of EFGC branch in our image2product search.
group feature fields intention of feature interactions
@@
It is crucial to organize feature fields into distinct groups and perform feature crossing within these specific groups, driven by domain knowledge. To achieve this, we can follow three main steps: 1) Identify Desired Matching Patterns: Determine the types of matching patterns that need to be emphasized based on the problem domain. 2) Group Feature Fields: Select and group the feature fields in a way that aligns with the identified matching patterns. 3) Integrate Group Embeddings into Subsequent Layers: Input the embeddings from each group into the following layers to facilitate the learning of feature interactions. For instance, in step 1 of our image2product search, we aim to emphasize the visual relevance between query image and item image, while personalizing the results based on each user’s preferences on item attributes. Step 2 is formulated as:
@@
Remark: Through splitting feature fields into groups and performing feature crossing within these groups, EFGC branch enhances the model’s ability to memorize specific patterns and avoid undesired feature crossing. Additionally, it is extensible to include more types of feature groups as long as desired matching patterns are given. The proposed EFGC is extensible and easily applied to other fields by using the designed feature groups in Table 1.
We note feature engineering techniques also use specific feature fields to create new features, such as the count of clicks between userID and brand. Both feature engineering and EFGC employ domain-driven knowledge. However, the difference is feature engineering creates raw input features, whereas EFGC learns feature crossing within designed groups, enabling the automatic discovery of patterns among feature fields within groups. They are not mutually exclusive and can be used together in practice.
@@ 3.3. Deep Net and Low Rank Cross Net
where $\mathbb{FC}$ denotes one MLP layer for dimension reduction. $f^{\text{Cross}}$ denotes a mixture of low-rank cross network blocks. Due to the page limit, we do not elaborate it in detail here. For a detailed formulation, please refer to Appendix A or (Wang et al., 2021). After the feature interaction learning of each branch, we further include a shared top layer $f_{\theta}(\cdot)$ to reduce parameters and regularize training. The output latent features of different branches are formulated as:
@@
3.4. Multi-branch Cooperation
@@ 3.4.1. Branch Co-teaching
Different branches of MBCnet possess distinct advantages and inherently focus on modeling different patterns. These patterns, in turn, are represented by the training samples. As a result, the learning capabilities of different branches can vary significantly across different samples. On particular samples, if a robust branch has been effectively trained, it can assist a comparatively weak branch in enhancing its performance. We formulate this collaboration idea as ”branch co-teaching” objective function, whereby given specific samples, the strong branch provides guidance to the weak branch in training.
@@
where $BCE$ means binary cross entropy loss. Since CTR prediction is usually regarded as a binary classification task, we employ $-\log(0.5)$ as a threshold to assess the learning progress of each branch [^3] The prediction probabilities may be affected by the data distribution, we experimented with various thresholds and found $-\log(0.5)$ generally did the best.. When the loss value is less than $-\log(0.5)$, we consider the branch on this samples is learned effectively. Otherwise, we consider the branch is not learned well.
Branch Co-teaching Objective Formulation: Given the above defined strong and weak branches, we take the prediction of strong branches as soft label to guide the learning of weak branches on selected samples. The co-teaching loss is written as follows:
@@
where $i,j\in\{\text{EFGC},\text{Deep},\text{Cross}\}$. SG means stop gradient. $I^{ij}\in\{0,1\}$ is the indicator that only equals 1 when $\boldsymbol{p^{i}}$ is strong and $\boldsymbol{p^{j}}$ is weak shown in Eq. 11. Similarly, $I^{ji}\in\{0,1\}$ is the indicator that equals 1 only when $\boldsymbol{p^{j}}$ is strong and $\boldsymbol{p^{i}}$ is weak shown in Eq. 10. In other cases, both $I^{ij}$ and $I^{ji}$ equal 0, which means we will not perform co-teaching optimization on those samples. This suggests that when two learners do not produce significantly different predictions, transferring teaching guidance may be unnecessary. It allows both learners the flexibility to discover and capture their own patterns. We conducted an experiment to demonstrate this idea in Section 4.4. $C$ is the number of non-zero values calculated from $I^{ij}$ and $I^{ji}$, indicating how many disagreement pairs are used in $\mathcal{L}_{BCT}$.
Remark: The main idea of this bidirectional branch co-teaching objective is that the teaching guidance is only conducted on selected samples where two branches show disagreed predictions. It is one important difference between this approach and the knowledge distillation in (Zhu et al., 2020) which performs full sample supervision between teachers and students. Given specific samples, learning abilities of different branches vary and are gradually improved, which suggests that the branches cannot offer appropriate supervision for all samples, especially in the early training stage. Employing supervision on all samples could easily introduce noisy supervision, and this is why EnKD (Zhu et al., 2020) uses pre-trained teachers and temperature scaling tricks to mitigate the effects of such noise. In contrast, our branch co-teaching operates in an end-to-end manner, focusing solely on samples with disagreeing predictions to minimize supervision noise.
@@ 3.4.2. Moderate Differentiation
In MBCnet, multiple feature interaction branches make predictions for the same input sample. When these branches are supervised on the same sample, they have risk to learn identical feature interaction patterns. This redundancy can significantly hinder the model’s capacity to explore a variety of feature interactions. To enhance the model’s robustness, it is better to enable distinct branches to learn differentiated latent features, encouraging them to capture diverse interaction patterns even on the same sample. However, excessive differentiation between these latent features may disrupt the consistency of shared patterns across branches, potentially harming model performance. In this context, we propose ”moderate differentiation” for the latent features $\boldsymbol{z^{i}}$ across branches. This approach promotes the exploration of diverse interaction patterns while ensuring a balance between differentiation and consistency.
@@
where $W^{ji},W^{ij}\in\mathbb{R}^{d\times d}$ are two learnable matrices and they are inverses of each other. To avoid feature collapse where multiple $\boldsymbol{z^{i}}$ collapse to one point after transformation, we employ orthogonal transformation, a specification of equivalent transformation. This means $W^{ji}$ and $W^{ij}$ are the transpose to each other, i.e., $W^{ij}=(W^{ji})^{T}$. It is important to point out that the above constraint is not equivalent to $\boldsymbol{z^{i}}$ and $\boldsymbol{z^{j}}$ are orthogonal. The orthogonal transformation is a special case of equivalent transformation, and enables $\boldsymbol{z^{i}}$ and $\boldsymbol{z^{j}}$ to maintain a moderate-level differentiation. It has an advantage of preserving inner product of vectors, namely it can keep similarities of vectors after transformation.
Moderate Differentiation Regularization: Taking the above analysis into consideration, we can write the regularization loss for moderate differentiation as:
@@ 3.4.3. Branch Fusion and Training Objective Function
where $\alpha$ and $\beta$ are two hyper-parameters to weight the importance of loss terms. Notice that we use an arbitrary sample to succinctly demonstrate above objective function with simple symbols. When using the objective function in practice, we compute the expectation of loss values of mini-batch samples during training.
Discussion: Popular CTR works (Cheng et al., 2016; Wang et al., 2021; Guo et al., 2017; Mao et al., 2023) investigate various feature interaction techniques for improved performance. However, each technique offers its own advantages, and relying on just one may not adequately capture the patterns in complex datasets, especially in industrial scenarios with enormous users and items. Meanwhile, these methods usually simply combine them with prediction voting or latent feature concatenation. By contrast, the proposed MBCnet advocates different branches to explicitly learn from each other and cooperate together to form a more powerful model, which is a new idea in CTR ranking. Compared to recent research (Zhu et al., 2020; Liu et al., 2024) that also leverages supervision from other models, MBCnet implements a distinct and more advanced learning framework. The comparison results presented in the following sections will highlight this difference.
We have also made some theoretical understanding of our cooperation scheme. In ensemble learning (Zhou, 2025) and co-training (Han et al., 2018), usually two or more learners with different parameter initialization are used to keep learner divergence. This divergence helps the whole model to filter different types of errors and improve the model generalization ability. However, these learners would gradually converge to be close to each other as the training epochs increase because of rough learner alignment, and lose learner divergence for generalization. The cooperation scheme in MBCnet provides an explicit and effective way to keep the divergence in training. First, the branch co-teaching transfers supervision knowledge from strong branches to weak branches on disagreement samples. In other words, within the whole training epochs, two branches can learn from each other and always select disagreement data for training, so the divergences of branches will be always maintained. This can also be supported by the visualization in Figure 4 (d). Second, the moderate differentiation provides a good way to ensure learners to have a moderate-level learning divergence, and avoids both too large or too small divergence for this CTR learning.
@@ 4.1.1. Datasets
Since there are no public datasets with hundreds of rich feature fields, we conduct experiments on two industrial datasets: Pailitao-12month and Pailitao-24month. Both are collected from our image2product search, also known as “Pailitao” at Taobao. Pailitao-12month and Pailitao-24month contain user search behaviors ranging from 2023-06-01 to 2024-05-31 and 2022-06-01 to 2024-05-31, respectively. For both datasets, there are various feature fields, including query image feature, user profiles, user statistical features, item image feature, item attributes, item statistical features and some other designed clustering and crossing features. The dataset statistics are given in Table 2, which show good resource to verify our method for the industrial track submission. All experimental models utilize the same input features to ensure a fair comparison of their modeling differences.
@@ 4.1.2. Baselines
We leverage different strong algorithms for comparison including single branch methods and branch ensemble methods. The single branch methods are as follows. 1) DNN is a deep multi-layer perception model, 2) CrossNetV2 (Wang et al., 2021) is a low-rank version of cross net (Wang et al., 2017a) to explicitly learn high-order feature interactions. The branch ensemble methods are as follows. 3) Wide&Deep (Cheng et al., 2016) combines a wide set of cross-product feature transformations and deep neural networks. 4) DCNv2 (Wang et al., 2021) contains mixture-of-expert based low-rank feature crossing and DNN to learn bounded-degree feature interactions. 5) MMOE+ (Ma et al., 2018) originally is a multi-task learning model. We employ its mixture-of-expert encoder to work as a baseline with multiple-expert ensemble. 6) EnKD (Zhu et al., 2020) studies knowledge distillation technique for model ensemble. 7) FinalMLP (Mao et al., 2023) designs feature gating and interaction aggregation layers for an enhanced CTR model.
Table 3. The AUC and LogLoss comparison results of different models on our two datasets. Best results are in bold and the most competitive public baseline results are underlined.
@@ 4.1.3. Parameter Settings
In our experiments, we split the datasets by chronological order, with the last seven-day data as validation and test set, and the rest is taken as the train set. For all methods, we set the latent dimension of last feature learning layer as 128. The validation performance is set as early-stop condition in training. In MBCnet, $\alpha$ and $\beta$ both equal 0.1 by hyper-parameter searching. In EFGC branch, hidden units of each group crossing are [1024,128]. In CrossNet branch, the number of experts is 2 and each expert has 2 layers with low rank dimension as 16. The hidden units of dimension reduction layer in EFGC and CrossNet are both 512. In Deep branch, the hidden units are [2048,1024,512,512,512]. Hidden units of the shared top layer are [512,256,128]. Hyper-parameters of baselines are searched to ensure each achieves its best performance. Core codes are available in the Appendix.
@@ 4.2.1. Offline Comparison
We make model comparisons with recent competitive baselines on our two large-scale industrial datasets. The results are given in Table 3. From this table, we can observe that: 1) our MBCnet achieves obvious improvement over recent competitive baselines. It has a 0.61% and 0.72% AUC increase over the most competitive model on two datasets, which are impressive improvements in large-scale industrial data scenarios. 2) Even when solely using our EFGC model, it can achieve superior model performance over other baselines, demonstrating its effectiveness. EFGC is a simple but effective design, and can be regarded as a variant of DNN branch. It has advanced capabilities for reducing redundant feature interactions and effectively learning domain-driven knowledge. 3) On Pailitao-24month dataset, single DNN outperforms other baselines. This trend is commonly observed in large-scale industrial settings, where the enormous data can overwhelm many modeling techniques. Even in this context, MBCnet still achieves superior improvements, demonstrating the effectiveness of our approach that leverages cooperative multi-branches to predict user behaviors.